UT 20230922: ECD - LBTI and AO restart



Night Info

For the night of: Fri, 2023 Sep 22 ---> Sat, 2023 Sep 23
Local midnight = 2023 Sep 23,  7 hr UT, or JD 2460210.792
Local Mean Sidereal Time at midnight =  23 47 56.7

Sunset ( 1000 m horizon):   18 21 MST; Sunrise:   6 03 MST
Evening twilight:  19 39 MST;  LMST at evening twilight:  19 26
Morning twilight:   4 46 MST;  LMST at morning twilight:   4 35
12-degr twilight: 19 10 MST -->  5 15 MST; night center:   0 12 MST

Moonset :  23 22 MST   
Moon at civil midnight: illuminated fraction 0.546
0.5 days since first quarter, RA and dec:  18 24 00, -28 46.2

Weather:


Staff

  • OSA: Josh Williams
  • ISA/SO: Jenny Power
  • ECD Coordinator: Jenny Power
  • Special Guests: Jared Carlson, Steve Ertel

Summary

Night Plan

TCS 2023B update

TCS 2023B added the following features:
  • * The event system has been upgraded and now includes a throttling mechanism. You can now silence rapid and frequent events so they don't flood the TCS GUI. This mechanism is controlled by parameters in the "event.dat" file. So you will need to edit the file to use this feature. A technical note is included at: https://github.com/LBTO/tcs/blob/2023B/tcs/eventhandler/README. For now, if you wish to use this feature, please contact me or Paul for assistance. This feature also addresses IT#8377. This was contributed by Paul Grenz.
  • * The OSS subsystem has been updated to include new indicators for ARGOS arm status (locked or enabled), and other error status. You can also control (deploy and retract) ARGOS arms from the OSSGUI as well. Additional fixes have been added to address ARGOS arm movement issues from OSS. This feature also addresses IT #8916. This was also contributed by Paul Grenz.
  • * Dan Cox upgraded the ECS subsystem to support running chiller CH0402 in remote mode. However this feature requires some additional work on the telescope side. Dan will coordinate with the mountain crew to get it done in the next week or so. Until that happens, the ECS will still be running the chiller in local mode.
  • * Matthieu Bec fixed several low-level issues in the TCS that might cause unexpected system crashes. This will improve the overall robustness of the TCS system.
  • * Matthieu also added new reflective memory slots and data dictionary variable names for PEIPSI instruments.

The TCS 2023B is now available on the mountain, but it is not yet activated. Since tonight is an ECD night, it is highly recommended that OSA activates this new TCS before the start of the night so it can be checked out. To switch to TCS 2023B:

  1. ssh tcs@tcs1.mountain.lbto.org
  2. cd /lbt/tcs/current
  3. ./tcs-cluster switch 2023B
There is no need to interact with the commands, if everything runs okay, it will display "All commands completed successfully" in green on your terminal window.

If it runs smoothly, then please leave it running. If you encounter any issues with TCS 2023B, you can still revert it back to 2023A:

  1. ssh tcs@tcs1.mountain.lbto.org
  2. cd /lbt/tcs/current
  3. ./tcs-cluster switch 2023A

  1. ssh tcs@tcs2.mountain.lbto.org
  2. cd /lbt/tcs/current
  3. ./tcs-cluster switch 2023A
If you revert it back, then please report the issues to the software group.

  • OAC update implented

AZCAM Server update

For the new version of the AZCAM server, you can switch from the newer server (192.168.2.197) back to the old one (192.168.2.199). Currently, the new version is running.

  1. Log into tcs1: tcs@tcs1.mountain.lbto.org:
  2. Stop both GCS's: netconfig stop GCSL netconfig stop GCSR
  3. Check to make sure they are both stopped: netconfig ps
  4. Go into the directory where the config files for the GCS are: cd /lbt/tcs/current/gcs/configuration
  5. Move the working GCS config to a distinct name: mv GCS GCS_new_azcamserver
  6. Move the original directory back: mv GCS_working GCS
  7. Now, restart both GCS's: netconfig start GCSR netconfig start GCSL
  8. Verify they are now running: netconfig ps

LBTI AO and Science Cam
  • LBTI (2 x 0.5 night - need not be consecutive)
      1. AO functional check-out on sky, close loop on a bright star, bayside stage positions for having the star on the science camera as well. 2h, both sides in parallel. Any Conditions.
      2. Science camera and operations check-out in parallel to above, plus 1h when the AO is done. Any Conditions.
      3. Find focus on both sides. 30 min per side. Requires good conditions
      4. Play with NCPA to see if we can improve it. 1h plus any extra time if we’re faster on above tasks. Requires good conditions
      5. Find LMIRCam Fizeau fringes. 1h. Requires good conditions
      6. Try combination of M3 and roof mirror movement to rotate Fizeau fringes on LMIRCam, check/demonstrate that this has the intended effect. If time, try to align the fringes vertically on

SHARKNIR and LBTI parallel presets
Send a synchronous SHARKNIR and LBTI preset. Dependent on fix implemented daytime to issues seen UT 20230922

Pointing Model
  • Instructions for Pointing model here: https://wiki.lbto.org/Commissioning/ComTools#Pointing_Model
  • Use MODS for pointing model data collection.
  • PtStars lives in ServiceObserving /PointingModel. This does not seem to work properly on obs1 &2, so use 32bit machines or robs.
  • First start with a pointing, copointed, and collimation check on a star near PA +0 (on the positive side of the azimuth wrap).
  • Set up MODS with pointing mask in place, taking a test exposure to verify exp time (5sec?)
  • Set up pointing log through PCSGUI > Pointing Details > Pointing Log > Start.
  • Use ptmodel_collect to collect data. ptmodel_collect has two main modes, taking images with a science instrument or with the guide camera.
    • % cd data
      % mkdir -p 20131215/rdg
      % cd 20131215/rdg
      % idl
      IDL> iif_register, RDG
      % cd 20131215/ldg
      % idl
      IDL> iif_register, LDG
      *In each sided terminal
      IDL> ptmodel_collect

  • Use list generated from PtStars (PTStars_10m.dat as list for generation increments around in Az and el, update as needed)
  • Through PCSGUI, start Pointing Log
  • Sending active presets through IRAF or IDL, collimate (I'm more familiar with IDL presets)
  • Once collimated, take image
    • With MODS: use pointing mask, exp time set to make sure not saturated
    • If no science instrument (FBG) then no images to be taken, just collimate and move on

You need to take an image of the pointing star one way or another (science instrument or AGw). For the AGw the guider images are OK, but since so many are taken it is a big help to note what image corresponds to what pointing star (or copy them off to a directory somewhere).

You do want the telescopes co-pointed and collimates at the time you take the pointing model data. The co-pointing should be handled by the preset for MODS+MODS. Just wait until collimated if the guide star looks obviously bad. Collimation should be pointing-free, but it is harder to centroid on a blobby star.

FBG focal station checkout
  1. AGW1/2 Alignment
  2. M3 Alignment LFBG,RFBG
  3. Field Aberration

Closed Dome Plan


Night log

-- JenniferPower - 19 Sep 2023
Topic revision: r2 - 22 Sep 2023, JenniferPower
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback